(-3, -5); m = -2 in point slope form

Found 2 solutions by MowMow, ikleyn:
Answer by MowMow(42)

y - y1 = m(x - x1)
y + 5 = -2(x + 3)
y + 5 = -2x -6
y = -2x-11

Answer by ikleyn(53765)

.

When a slope "m" and a point (a,b) are given, then the point-slope form equation is

    y - b = m*(x-a).


In your case, the point-slope equation is 

    y + 5 = (-2)*(x+3).      ANSWER


This and only this form.  It is your desired equation.


Do not accept any other answer.

Solved.

-------------

This problem and this question do belong to the category of problems and questions, where you should not think:
you simply MUST know the answer.

Same as you know the multiplication table.

-------------

It is very elementary and basic skill that every student must develop to the level to solve such problems on his  (or her)
own  AUTOMATICALLY  without asking assistance from outside.

To help you to develop such skills,  a special lesson was created in this site

    - Equation for a straight line having a given slope and passing through a given point

Find there a complete explanation on how to do it and learn the subject once and for all.

Do not miss other associated lessons that cover more wide subject.

You will find the links at the end of the referred lesson.

Consider these lessons as your textbook,  handbook,  tutorials and  (free of charge)  home teacher.